OpenStudy (anonymous):

please help! x+4y-z=1 3x-y+8z=0 x+4y-z=10

OpenStudy (anonymous):

what is the value of each variable?

OpenStudy (anonymous):

pick two equations out of the three first so lets do the first two

OpenStudy (anonymous):

ok i did the question and i got 0=9 but i dont know what to do next!

OpenStudy (anonymous):

how did u do the question and how did u get 0=9

OpenStudy (anonymous):

pick x+4y-z=1 and 3x-y+8z=0

OpenStudy (anonymous):

i did the process of elimination and got rid of the x variable then after doing that on two of the equations i got 0=9 on both but i dont know what to do next!

OpenStudy (anonymous):

multiply 4 to 3x-y+8z=0

OpenStudy (anonymous):

lets cancel the y's first

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so did u multiply the 4?

OpenStudy (anonymous):

yup! i got 12x-4y+32z=4

OpenStudy (anonymous):

actually the 4 at the end is a 0 because 4*0=0

OpenStudy (anonymous):

oh whoops...

OpenStudy (anonymous):

12x-4y+32z=0

OpenStudy (anonymous):

yes then add or subtract the two equations now so x+4y -z =1 +12x-4y+32z=0

OpenStudy (anonymous):

so the y's cancel out and ur left with 13x+31z=1

OpenStudy (anonymous):

yea ok

OpenStudy (anonymous):

do you now do the same ot the other equation? cancel the ys

OpenStudy (anonymous):

yes u do, but actually this problem doesn't have an answer

OpenStudy (anonymous):

it doesnt? so its like a no solution vertical line?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

sorry, i should have figured that out earlier, i made u go through all this mess and wasted ur time......

OpenStudy (anonymous):

no its ok :)

OpenStudy (anonymous):

well u got it now :)

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

ur welcome :)